NvidiaIntely (nvidiaintely):

In a nuclear reaction, the mass defect value = 0.0304 kilograms. What amount of energy, in joules, does this loss in mass represent? (1 J = 1 kg m2/s2) 2.74 x 1015 J 9.00 x 1018 J 9.00 x 1010 J 6.04 x 1014 J

OpenStudy (will.h):

Gosh this one looks awful lol

OpenStudy (will.h):

E= 0.0304*(3.0*10^8)^2 it is A 2.74 x 1015 J Gotta go have dance class byeee

NvidiaIntely (nvidiaintely):

Thanks, Bye WIll!

OpenStudy (kevin):

0.0304 x (3 x 10^8)^2 = 2736000000000000 (A)

NvidiaIntely (nvidiaintely):

yeah

OpenStudy (kevin):

lol